OpenStudy (anonymous):

R=1+2cos(theta) " find all theta when r=0 ; 0

OpenStudy (mertsj):

\[2 \cos \theta =-1\] \[\cos \theta =\frac{-1}{2}\] cosine is negative in quadrants II and III